OpenStudy (avanti):

Evaluate cos(Sin-1 0) Is it 1?

OpenStudy (jdoe0001):

\(\bf cos(sin^{-1}(0)) \ \ ?\)

OpenStudy (jdoe0001):

if that's what you're asking, yes, is 1

OpenStudy (avanti):

This is what the question looks like.

OpenStudy (jdoe0001):

hmm, then yes, sine is 0 at angles \(\bf \cfrac{\pi}{2} + n\pi; n >0\) at those points, that is 0 and pi for one revolution, then cosine will be \(\bf \pm 1\)

OpenStudy (avanti):

thank you so much!

OpenStudy (jdoe0001):

yw

OpenStudy (jdoe0001):

hmm, actually anyhow, should have been \(\bf n\pi; n >0\)

OpenStudy (jdoe0001):

well, \(\bf n\pi; n \ge 0\) better

OpenStudy (jdoe0001):

hehe, sine is 0 at pi and 0 and further the same on whatever revolutions